# From the CDC to emerging infectious disease publics: The long-now of polarizing and complex health crises

### v2

@@ 8. Discussion
At the top of this paper, we introduced the disaster of the smallpox outbreak in Milwaukee 130 years ago. Then, as now, public health crisis communication is quite political. The main difference that current-day public health institutions have to grapple with is a more complex sociotechnical communication ecology.
We argue that platform affordances and emerging technologies, especially in generative AI, are all important to incorporate into the public policy discourse framing analytical framework (Figure 1). We first discuss the lack of crisis message clarity in a continuously evolving pandemic enmeshed in polarized political framing seen through the lens of Twitter affordances in §8.1. Next, we propose design recommendations based on our findings, specifically, a CDC AI assistant in §8.2. Finally, we present policy implications for our findings to support the CDC and other public health organizations in managing online communication for emerging infectious diseases in §8.3.
8.1. Crisis messaging journeys, platform affordances, and verified users
In this section, we reflect on our research questions to better understand CDC communication through the lens of Twitter affordances and message properties (§8.1.1), polarized publics (§8.1.2), and message journeys through pandemic flashpoints (§8.1.3).
8.1.1. RQ1: Engagements, sentiments, and Twitter affordances
The importance and low levels of two-way communication
There were more than 21 thousand unique Tweets responding to the CDC directly, averaging five retweets for each Tweet, using rich media like “receipts” which included earlier messages from the CDC (see §5.1.2). However, the CDC itself only responded directly 58 times to 25 unique accounts. This result is in line with earlier findings about how the CDC and related organizations used Twitter as top-down communication channel in H1N1 (Gesser-Edelsburg et al., 2014) and Ebola (Dalrymple et al., 2016). Two-way communication is central to the CERC model (Reynolds and Seeger, 2005) because it increases mutual respect and the potential to incorporate the views of the public into organizational communication practices (Dozier et al., 2013). We propose design recommendations to support more bottom-up communication in the future in §8.2.1. Public understanding of science is made in collaborative communication between scientists and members of the public (Felt and Davies, 2020). However, this communication is especially susceptible to being taken out of context and/or used in misinformation (Allcott et al., 2019) which we discuss next.
Low-credibility sources, verified accounts, and rich media
We found that users who use low-credibility sources are more likely to continue posting on the CDC public and to mention other users in the network, while using the rich media which can increase the persuasiveness of misinformation (Zhou et al., 2021a). While low-credibility Tweets themselves might not be propagated at a high rate, they might be when created and shared by verified accounts. In fact, as we described in §5.1.3, we found that one of the verified journalists (@alexberenson) has used less credible sources in a number of his Tweets which countered the CDC’s mitigation messages (e.g., lock-downs, masking, and vaccinations).
Our findings show that the use of rich media increases the propagation of Tweets (see Table 1) and that Tweets with low-credibility sources were more likely to use rich media (see 5.2.5). Therefore, the development of more complex generative text2image and text2video models (Zhao et al., 2023) which allow users to quickly produce rich media, presents another future moderation challenge. While Tweets using less credible sources were not heavily propagated in our dataset (see §5.1.1), it is difficult to determine if this is caused by any platform-level algorithmic moderation intervention (c.f., (Duskin, 2024; Sharevski et al., 2022)). Similar content, though not using the same sources, could be propagated at higher rates. This calls for more flexible and adaptive moderation models that could find and counter this content. We describe one in §8.2.2.
Sentiment valence not always representative of message framing
While earlier work took negative sentiment to be a proxy for opposing public health policy (e.g., (Unlu et al., 2023)), our work shows that negative sentiment might indicate that the user does not think the public policy went far enough (e.g., reducing the number of quarantine days; see §5.2.3). Similarly, we described in §5.2.2 how posts with neutral sentiments can provide nuanced critiques of current policy while also forwarding public health messages in local publics (e.g., high-school event guidelines). We continue our analysis of affective publics (Valiavska and Smith-Frigerio, 2022) next.
8.1.2. RQ2: Polarized affective publics and public health
In §6, we presented a number of publics along with examples of affectively polarized frames (Tables 9, 12, 16, and 17) along with an array of stakeholders (e.g., teachers, public health professionals, parents, racial equity advocates) who in many cases espouse an us vs. them engagement (i.e., #MAGA vs. #VoteBlueNoMatterWho) while framing central issues like vaccination in two different planes - public safety vs. personal freedom (i.e., #vaccinessavelives vs. #vaccinerisk) (Brüggemann and Meyer, 2023).
This polarization was also evident in our network analysis. We found that the high likelihood of reciprocity coupled with the sparsity of the mentions network centered around the CDC (§5.1.3) suggests the creation of filter bubbles (Barta and Andalibi, 2021). When analyzing propagation, we found that discussing polarizing issues like vaccinations and lock-downs increased the chance of retweets and likes which might also provide another avenue for the creation of filter bubbles (see §5.1.1).
All these findings suggest that users on the CDC public see filtered content based on who they follow, the hashtags they search for, and suggestions made by recommendation systems (filtering algorithms). In turn, the behaviors place users in bubbles of curated information (i.e., filter bubbles;(Barta and Andalibi, 2021; Sundar and Lee, 2022). While earlier work shows that filter bubbles can polarize people by distorting their realities in their echo chambers (Pariser, 2011), marginalized communities can benefit from being in such filter bubbles where they can find other users with similar experiences (Randazzo and Ammari, 2023). For example, cluster 1, which focuses on racial health disparities of COVID on marginalized communities, might be a beneficial space for organizations (e.g., ADOS), researchers, and others who might be interested in this issue. If a new entrant to the network mentions the CDC, or uses #CDC, and asks about racial disparities in health outcomes in the context of COVID, then the recommender system can introduce them to central nodes in cluster 1. We describe a relevant design recommendation in §8.2.4.
Given its mission as a public health institution, the CDC needs to reach out to people “where they are” (Mathews et al., 2020), we will introduce a design recommendation for reaching out to different issue publics in §8.2.4. We argue that by finding and engaging with publics of interest, the CDC can focus on crafting narrative messages that can fit the needs of each community (Sandell et al., 2013; Alamir et al., 2024).
8.1.3. RQ3: Managing friction in crisis messaging journeys from one flashpoint to the next
In §2.3.3, we discussed how it was difficult to gauge different phases of a public health crisis, especially when it is an emerging infectious disease, like COVID. Dalrymple (Dalrymple et al., 2016) argues that there is a need for a rapid change of communication strategy around these flashpoints, and that, due to social media affordances (§2.1), especially persistence and forwardability, changing communication strategies can be difficult.
In Study III (§7), we used change-point detection and an analysis of Tweets before and after those events to describe two phases that hinged on a change in guidelines based on new data and a better understanding of the virus as well as the introduction of new factors like vaccines and new virus variants. In both of these cases, responses to the CDC included “receipts” of earlier CDC messaging. This might be explained by the perceived lack of honesty about the CDC’s uncertainty vis-a-vis an emerging infectious disease (Holmes, 2008). In fact, Holmes et al. (J. Holmes et al., 2009) suggested that focusing on “certainty for now” could earn public trust more than conveying certainty and then being proven wrong about IEDs in future flashpoints.
This is in line with earlier work that shows how data friction (challenges of sharing and interpreting data) (Edwards et al., 2011) increases with the distance between message creators and re-users (Borgman and Groth, 2025) and with the duration of time from the original message to the time of reuse (Bates et al., 2016; Borgman and Groth, 2025). The different “ways in which [data] are …modified to fit different uses as they travel across space, time and social situations” are defined as data journeys (Leonelli, 2020). We introduce the concept of crisis messaging journeys which describes how crisis messages initially shared by the CDC change based on context, purpose (e.g., opposing new mitigation guidelines), and time elapsing from the initial CDC message (Gregory et al., 2024).
Guidelines shared by the CDC at any time during the pandemic might change later given new information gleaned about the EID (e.g., while the vaccine is still effective against hospitalization, people can still be infected with new variants). While scientists understand that “scientific knowledge is always provisional, always open to revision,” (Edwards, 2013, P.427) public engagement with scientific knowledge (i.e., members of the public using “receipts” of earlier guidelines to argue against the veracity of updated guidelines) can “represent ideological and political strategies [that] represent a new form of friction…inclusion friction [which can] promote confusion, suspicion,” and misinformation (Edwards, 2013, ibid).
There are three ways to reduce inclusion friction: (1) using more cautious language (e.g., employing caveats in public science communication (Stocking and Holstein, 1993, P.200)) when presenting guidelines signifying “this is what we know now, rather than this is what we know”(Gregory et al., 2024); (2) linking the original message either by quote-tweeting it to explicitly show the evolution of science and/or annotating (Zheng and Ma, 2022) the original message using an affordance like community notes (Kankham and and, 2024); and (3) pre-exposure (Lewandowsky et al., 2012, P.116) or pre-bunking (Barman and Conlan, 2024) where the audience is presented with an alternative narrative (Lewandowsky et al., 2012, P.117), thus filling any gaps which might result from changing messages in a flashpoint. This would include messaging to explain the reasons for new guidelines by comparing them with their original messages and explaining how the difference was brought about by new medical information. A mechanism to detect flashpoints in real-time is presented in §8.2.3.
Earlier work has shown that if uncertainty claims are made early by scientists in their publics science communication, they are seen as being frank as they “openly admit incomplete knowledge.” (Zehr, 2012, P.9) This is imperative to the success of the crisis organization because the publics’ audiences are more “likely to reject organizational messages during a crisis situation that are inconsistent with past actions” (Stephens et al., 2005, P.399) which would negatively effect crisis communication around flashpoints.
8.2. Design Recommendation: The CDC AI Assistant
Sanders et al.(Sanders et al., 2022) introduce a generative response model based on earlier public health messaging. Indeed, while social media interactions are usually between human actors, new generative AI models like ChatGPT are introducing non-human agents (Quéré et al., 2025; Aubin Le Quéré et al., 2024; Schroeder et al., 2024) which could play a bigger role in exchange of health-related information. Endacott and Leonardi (Endacott and Leonardi, 2022) build on this idea through the conceptualization of artificially intelligent communication technologies (AICTs), defined as “goal-directed, computational agents that use AI to make decisions about communication and communicate on someone’s behalf.” New Generative AI models introduced on social media sites like Facebook are interacting with users of the platform. Meta AI on Facebook, for example, can respond to questions through messenger (Meta, 2025). This is an example of advanced AICTs, or Generative assistants (Park et al., 2023) which have an “architecture that handles retrieval where past experience is dynamically updated at each step and mixed [the] current context.” Extending the design of current offerings (i.e., Meta AI), in this section, we present the CDC AI assistant, which is designed to support the CDC staff based on three triggers: (1) direct response to the @CDCGov account; (2) suspected mis/disinformation (a flag that could be raised by other users or using automated moderation tools); and (3) temporal change-point detection.
Though they can perform helpful tasks, introducing advanced AICTs reduces the organization’s “degree of control over the communication environment” (Park et al., 2023). That is why we need a human-in-the-loop to work on verification of triggers (e.g., is this actually misinformation) and, where necessary, editing of messages to be published by the CDC AI assistant (Wu et al., 2022).
8.2.1. Tailored responses to responses tagging the CDC account
The CDC AI assistant can be trained on earlier messages shared by the CDC account on Twitter as well as content produced on other online accounts. After fine-tuning a generative AI on earlier CDC text, it can reproduce information shared earlier by the CDC on Twitter or available on the CDC website. This can be done in response to the first trigger when the CDC is responded to directly. However, given the sensitivity of messages from the CDC, we will model the assistant after the CoAuthor AI assistant developed in (Lee et al., 2022) where the assistant provided a sounding board to authors. More specifically, given the text of the response to the CDC account, the assistant can present the CDC staff (i.e., the human-in-the-loop) with suggestions based on earlier text and their context (Zhang et al., 2023). To help the CDC staff further, the AI assistant can explain the suggestions using chain of thought (CoT) represented by a step-by-step breakdown of the reasoning (Wei et al., 2022). After the human-in-the-loop drafts the final message, they can save it to an archive of responses along with the original response to the CDC and any annotations they might find helpful. These responses can be used to retrain the AI for future engagements.
8.2.2. Detecting and responding to dis/misinformation
While social media platforms can introduce system-level moderation tools to detect misinformation, these applications are usually not trusted by social media users who would rather use tools that are produced by organizations that do not have a profit motive (Jhaver et al., 2023). Additionally, as crises evolve, the nature of misinformation evolves quickly as exemplified by the COVID “infodemic.” (c.f.,(Wang et al., 2023; Hobbs et al., 2024)) That is why relying on the wisdom of “crowds of ordinary users” to flag dis/misinformation by accounting for “content facticity, user intent, and perceived harm” has proven to be a more effective moderation policy (Bozarth et al., 2023). In our case, users might trigger the CDC AI assistant by mentioning @CDCGov and using a hashtag like mis/disinformation that would then again be annotated and responded to by CDC staff as introduced above. As a substantial repertoire of mis/disinformation content is built up, the assistant can be trained to respond to messages that repeat misinformation even if worded differently by searching for similar text. In addition to community moderation flags, the assistant can also be trained on annotated mis/disinformation datasets (c.f., (Hayawi et al., 2022)) or by flagging low-credibility sources as we did in §2.1. This would allow the assistant to find and counter more disinformation (Sahu et al., 2021).
8.2.3. Real-time flashpoint detection
We analyzed change-points based on content in §7.1 and based on message content and user properties in §7.2. Given the importance of identifying flashpoints and changing communication strategies accordingly (Dalrymple et al., 2016), automating the detection of change-points can help the CDC deal with flashpoints as they occur. Based on a system presented in (Lyu et al., 2024), we first apply a baseline change-point data set from earlier interactions (as presented in §7), which are then incorporated into a Bayesian model. This first phase provided one of two outputs: either the change is a candidate for a change-point or it is not. If it is, a PELT model is applied at set time durations (e.g., every week or month). When detecting a change-point, manual inspection is still important to verify that indeed there is a change after the flashpoint. The manual coder can confirm that indeed, this is a flashpoint, remove it, or if not clear, set it as pending to request annotation by another human-in-the loop. This will allow the CDC to change its communication strategy.
8.2.4. Engaging with different publics and helping users find communities
Given how we described the importance of providing different narratives for different polarized publics (see §8.1.2), creating a multi-agent system would allow the CDC to engage in a more sustained communication process with more users by allowing different generative agents to focus on different publics (Hadfi and Ito, 2022) represented by the clusters we identified in §6. For example, we would build a CDC AI assistant focused on racial bias in healthcare. That agent would engage most with the community in cluster 1 by providing updates in relation to racial bias (e.g., new data released) or responding to messages in the cluster if they are not answered in a specified amount of time (e.g., 12 hours).
The assistant can also suggest accounts at the center of cluster 1 for users interested in the primary topics discussed in the cluster especially if they have not engaged with them before. Given that we are using generative AI, the assistant can also explain why they are being suggested user(s) based on their interests (Lubos, 2024; Di Palma, 2023). For example, the assistant could message a user saying “given your interest in healthcare equity, you might want to follow [list of users in cluster 1].”
8.3. Policy Recommendation: Preparing for the long-now of health crises
Most studies of crisis informatics focus on the response of crisis organization (e.g., the CDC) during early phases of a crisis (Soden and Palen, 2018). However, even in early 2025, on the eve of the fifth anniversary of COVID, around 40% of the population still think the pandemic is “not over,” that they are not back to living their “normal” lives, and are worried about the long-lasting effects of COVID (e.g., long COVID) and of the next health crises (e.g., bird flu) (Gallup, 2025). This shows that there is a need to understand the long-term communication response needs for public health crises like COVID.
Continued engagement with issue publics, especially those with well-defined stakeholders and constituencies (e.g., racial equity and pain advocates), and regional publics (e.g., local help groups providing mutual aid (Soden and Owen, 2021; Randazzo et al., 2025)) can prove especially important in the future should there be any engagement with local populations and their constituencies. Earlier work on local mutual aid groups showed that while they started by addressing immediate needs (e.g., access to food) resulting from the early COVID surge, these same groups advocated for rent control a few months into the crisis (similar to the rent control advocates identified in §6.2.2). Additionally, as mitigation guidelines started being relaxed, the needs of immunocompromised, persons with special needs (Zhang et al., 2022a; Joy et al., 2024; Mack et al., 2022a), chronically ill (Mack et al., 2022b), and long COVID (Karra et al., 2023) populations become paramount.
#CDCChats like the ones presented in (Dalrymple et al., 2016) can be a start, however, the CDC might want to use different platforms depending on population needs. For example, recent work found that many uncertain parents relied on WhatsApp connections to make better sense of COVID (Joy et al., 2024). The design recommendation presented in §8.2 can support the CDC with long-term engagements with different stakeholders. Additionally, the use of data annotated by the human-in-the-loop described above can create more contextualized data. In doing so, Twitter user engagements can provide citizen sensors (Goodchild, 2007) where citizens can “create and share information, forming vast sensing that allow for information in certain topics to be collected, stored, mined, and analyzed in a rapid manner.” (Huang et al., 2022, P.2).
Since earlier work shows how online and offline publics overlap in their crisis concerns (Randazzo et al., 2025), the CDC can model signals from specific regions of the country over time to determine policy changes in real time. This can be done by combining methods introduced by Wen et al. (Wen et al., 2022) who discussed how news topics changed over time in response to policy changes and the work done in Whata & Chimedza (Whata and Chimedza, 2021) who used a dataset of government measures (e.g., lock-downs) and the Google COVID-19 Community Mobility Reports dataset which measures population density in several public space categories (e.g., pharmacies, parks, etc.) over time to determine the effects of government policies in different locales.
Reflecting on the public policy discourse framing analytical framework adapted from (Park and Lee, 2020) which we present in the introduction (Figure 1), we introduce three new components to the model, which we argue should be considered by crisis-management organizations as they engage with the public:
  1. (1)
    Social media affordances: what messages are propagated most and how networks of interactions are formed.
  2. (2)
    Message and user properties: sentiment valence of the message, whether it uses low-credibility sources and/or rich media, and whether it is shared by verified Twitter accounts are all important to take into account in crisis communication.
  3. (3)
    The crisis message journey throughout the “long-now” of public health crises as new flashpoints resulting from new scientific findings (e.g., new variants) and changes on the ground (e.g., vaccine roll-out) are important to manage as they arise.
Additionally, we argue that it is important to understand the different issue, affective, and regional publics (e.g., pro-mitigation or racial health equity clusters). Finally, given the impact of generative AI on how people create content and engage with others, crisis organizations should consider employing these technologies to support their communication strategies while also developing measures to counter their effects on crisis communication.

### v3

@@ 6. Discussion
In this section, we reflect on the implications of our findings for understanding how crisis discourse unfolds across fragmented digital publics. In §6.1, we reflect on how our findings relate to earlier work. In §6.2, we present design recommendations to better support both crisis organizations like the CDC and other publics. Finally, we provide policy recommendations for crisis organizations as they create and maintain long-term engagements with different publics (see §6.3).
6.1. Reflecting on Earlier Literature
To interpret the communication landscape that emerged around the CDC during the COVID-19 pandemic, we begin by unpacking the affective and ideological dynamics that shaped public engagement. In Section §6.1.1, we examine how partisan identity and political framing contributed to the formation of polarized discourse clusters that often operated in ideological isolation. In §6.1.2, we shift focus to reflecting on the spectrum of stakeholders and emotionality, exploring how affect—particularly anger, fear, and sarcasm—served both as a mobilizing force and a barrier to sustained cross-cluster dialogue. Finally, in §6.1.3, we analyze the structural configuration of discourse networks, highlighting how most clusters functioned as echo chambers while only a few enabled cross-cutting communication. Together, these subsections show how pandemic-related discourse was shaped not only by the content of messages but by the affective, ideological, and structural conditions under which they circulated.
6.1.1. Crisis Communication in Context: Reflecting on Frames and Affective Polarization
Our findings reinforce prior research on the centrality of framing in public health crisis communication, especially when mediated through social media platforms. The CDC public on Twitter emerged as a highly polarized, affect-laden environment where hashtags served as tools of alignment and boundary marking. As with earlier work on issue and affective publics (Papacharissi, 2016; Brüggemann and Meyer, 2023), users participated in identity-driven framing of core topics such as masking, vaccination, and school reopening. These frames often manifested in polarized pairings: public health versus personal liberty, science versus politics, and equity versus neutrality. This duality was most starkly evident in how users interpreted CDC messaging—either as lifesaving guidance or authoritarian overreach.
Emotion, or affect, played a critical role in shaping these interpretations. Hashtags functioned as affective signposts, clustering users into emotionally resonant communities. For example, frames like #VaccinesSaveLives invoked collective responsibility, while #VaccineRisk or #Plandemic expressed fear and distrust. As found in prior work (Jin, 2010; Kim and Cameron, 2011), messages that connected with pre-existing values and affective identities gained the most traction. Thus, affect did not just shape response but served as a filter through which users interpreted public health information.
6.1.2. Stakeholders Across Publics: A Spectrum of Engagement
Our analysis of the CDC Twitter public revealed diverse stakeholder groups, which aligned closely with Hallahan’s typology (Hallahan, 2000). Active stakeholders—such as healthcare professionals, teachers, and advocacy organizations—were visible in clusters focused on mitigation strategies and racial health equity (e.g., Clusters 1 and 2). Aware and aroused stakeholders, such as parents concerned about school reopening or travel guideline changes, populated Clusters 12 and 13. Inactive or minimally engaged users were harder to trace but may exist outside the dominant echo chambers.
Significantly, our findings demonstrate that different publics emerged not only around ideological lines but also around topical and affective interests. Cluster 1, for example, focused on racial health disparities and exhibited the highest level of cross-cutting engagement (E-I Index of 1.0), unlike most other clusters that were ideologically siloed. This cluster illustrates how issue-based publics can maintain sustained engagement around structural concerns even when broader discourse trends shift. Importantly, this cluster reflected a high concentration of verified health equity advocates, academic institutions, and non-profit organizations, underscoring the value of trust-building and collective advocacy in marginalized communities.
6.1.3. Cluster 1 and the Evolving Needs of Marginalized Publics
Cluster 1 stood out in its consistency and cohesion across the timeline. While discourse around race declined in the broader CDC public early in the pandemic, Cluster 1 remained focused on racial disparities in healthcare, particularly as they intersected with COVID-19 outcomes. Our time-series analysis revealed flashpoints of intensified engagement tied to external socio-political events, including the George Floyd protests and CDC announcements about racism as a public health threat.
The thematic discourse within Cluster 1 moved from highlighting disproportionate mortality and hospitalization rates among BIPOC populations to demanding data justice and structural accountability. These shifts illustrate a critical feature of crisis communication: the evolution of public concerns over time. Cluster 1 encapsulates how marginalized publics respond to both the pandemic and the ways institutions manage it. Their discourse challenges symbolic gestures (e.g., declarations without follow-through) and calls for tangible structural reform. As such, Cluster 1 offers a model for understanding the layered, evolving needs of issue-specific publics, especially those historically underrepresented in public health dialogue.
However, while Cluster 1 continued its engagement with other CDC publics, and was overall one of the more retweeted clusters, it had little overall effect as the decline in overall focus on health inequity at the intersection of race and class. This finding expands on ealier work by Ahmed who found that certain bodies and their stories are made illegible in public discourse due to emotional economies in which certain narratives are more easily absorbed or circulated than others (Ahmed, 2013). In particular, affectively charged narratives that align with dominant sensibilities—such as hope, resilience, or redemption—are often privileged, while those that center on structural critique, pain, or systemic injustice are marginalized or dismissed. As a result, even when marginalized publics gain visibility through mechanisms like retweets or temporary virality, their ability to reshape dominant narratives remains constrained. This underscores the limits of visibility as a proxy for influence and highlights how affective economies continue to regulate which stories matter and which are rendered unintelligible. In the next sections, we present design and policy recommendations to address this issue.
6.2. Design Recommendation: Multi-Agent CDC AI Assistants for Diverse Publics
Building on our analysis, we recommend the development of a multi-agent generative AI system tailored to specific issue publics identified in the CDC Twitter discourse. These AI agents, modeled on clusters such as racial health equity (Cluster 1), school reopening (Cluster 15), and mask guideline clarity (Cluster 7), can deliver targeted, responsive messaging and facilitate two-way communication.
A racial health equity-focused agent, for instance, could proactively disseminate updates on vaccine equity, respond to queries about structural racism in healthcare, and amplify voices from trusted organizations. It can also provide context-aware recommendations (e.g., suggest following key equity advocates), improving both trust and reach. The CDC AI Assistant should operate with transparency, explaining the rationale behind each interaction (Lubos, 2024; Di Palma, 2023). A similar structure was suggested to suggest new content and social connections in foster care online communities (Ammari et al., 2025).
Importantly, such systems should be co-designed with public health experts and community members to ensure cultural competence, appropriateness, and ethical data use. This approach aligns with inclusive design principles and expands the CDC’s ability to meet different publics ”where they are” (Mathews et al., 2020).
6.3. Policy Recommendation: Planning for the Long-Now of Crisis Communication
Traditional crisis communication frameworks often prioritize immediate response. However, our study affirms the need for sustained, adaptive engagement over the long arc of a public health crisis. The CERC model (Reynolds and Seeger, 2005) emphasizes pre-crisis relationship-building with stakeholders. Our findings extend this principle by highlighting the importance of long-term monitoring and interaction with issue-based publics across platforms.
To operationalize this, the CDC and similar institutions should incorporate real-time modeling of regional discourse dynamics, aligning with work by Wen et al. (Wen et al., 2022) and Whata & Chimedza (Whata and Chimedza, 2021). These models can help detect emergent needs, track sentiment changes, and inform timely policy adjustments. Collaborating with local mutual aid networks and community leaders can further bridge the gap between national messaging and local realities (Soden and Owen, 2021; Randazzo et al., 2025).
Finally, public health institutions should prepare for the integration of emerging technologies into their communication infrastructure. This includes not only leveraging generative AI tools but also instituting safeguards against their misuse. As our revised public policy discourse framing framework (Figure 1) suggests, platform affordances, user behaviors, and algorithmic governance must be central to designing resilient communication ecosystems for the crises of tomorrow.